If 5.03 g of Ethylene Glycol are added to 96.6 g of water what is the freezing point of the solution
Answer #1

first calculate molality of solution

molality = (W /MW) (W / mass of solvent g)

molality = (5.03 / 62.07) ( 1000 / 96.6)

molality = 0.84 M

now

Tf = Kf x m

Kf = 1.86 0C / m

Tf = 1.86 x 0.84

Tf = 1.56

Tf = 0 - 1.56

Tf = - 1.56 0C